OpenStudy (anonymous):

divide using synthetic division and write a summary in fraction form: (x^3-5x^2+3x-2)/(x+1) The answer should come out to be x^2-6x+9+(-11/(x+1)) but i have no idea how to get there. Halp!

OpenStudy (anonymous):

if you can open this, it will show you, if not i can walk you through it www.mathportal.org/calculators/polynomials-solvers/synthetic-division-calculator.php?formId=0&combo1=2&combo2=1&val4=1&combo3=2&val3=5&combo4=1&val2=3&combo5=2&val1=2&combo6=1&val6=1&rb1=div&ch1=expl

OpenStudy (tkhunny):

Did you do the synthetic division?

OpenStudy (anonymous):

I did it on my own separately and then i also found this and thought it'd be easier

OpenStudy (tkhunny):

What were the results across the bottom? Was it 1 -6 9 -11?

OpenStudy (anonymous):

@tkhunny yea

OpenStudy (tkhunny):

Perfect. You did it very well and you are done. You just don't know it!!! That's good news. :-) Compare the results of the bottom line to the coefficients in the "answer". x^2-6x+9+(-11/(x+1)) 1, -6, 9, -11 Coincidence? No.

OpenStudy (anonymous):

ya, I got that but i dont know if they got the same thing XD

OpenStudy (tkhunny):

Who are these "they"? If you did the synthetic division correctly (and you did), and obtained 1, -6, 9, -11, then you have proven that: \(\dfrac{x^{3}-5x^{2}+3x-2}{x+1} = x^{2} - 6x + 9 - \dfrac{11}{x+1}\). I'm not seeing the problem.

OpenStudy (anonymous):

Yes but im not the one that asked the question lol

OpenStudy (tkhunny):

?? Well, no one can help you if you fail to provide pertinent facts. Good luck.